OpenStudy (anonymous):

Medal for Best answer Multiply (2+sqrt12)(3-sqrt3) a 8sqrt3+12 b 4sqrt3 c-4 d 12-8sqrt3

OpenStudy (rulnick):

(2)(3) -2sqrt3 +3sqrt12 - sqrt(3(12)) = 6 - 2sqrt3 + 3sqrt(4)sqrt(3) - sqrt(36) = 6 - 2sqrt3 + 6sqrt(3) - 6 = ... Can you get it from here?

OpenStudy (anonymous):

4sqrt3?

OpenStudy (rulnick):

Yes!

OpenStudy (anonymous):

Thank you!

OpenStudy (rulnick):

welcome